# Stochastic Dominance Constrained Optimization with S-shaped Utilities: Poor-Performance-Region Algorithm and Neural Network

### v1

@@ 7 Conclusion
We study a utility maximization problem under stochastic dominance constraints. Starting from an S-shaped utility function, we derive the explicit optimal solution without a liquidation boundary under first-order stochastic dominance (FSD) constraints. For the more challenging SSD problem, particularly under non-concavity, obtaining an explicit optimal solution is inherently difficult. Motivated by the structural properties of the analytical theorem in the concave case, we introduce a Poor-Performance-Region Algorithm (PPRA). This algorithm efficiently identifies the candidate structure that a potential optimal solution should satisfy and proves effective in the vast majority of cases. In addition, extensive numerical experiments illustrate how the algorithm operates, confirm its broad applicability, and demonstrate its capability to handle the few exceptional cases.
Building on the structural insights provided by the PPRA, we further recognize the potential of neural networks in tackling SSD problems under non-concavity. While neural networks offer strong approximation capabilities, their direct application often suffers from slow convergence and severe sensitivity to local optima induced by non-concavity. By leveraging the PPRA’s ability to capture the essential structure of the optimal solution, we develop an algorithm-guided piecewise-neural-network framework that integrates these structural cues into the learning process. Compared with a standard neural-network approach, this framework achieves significantly faster convergence and effectively avoids undesirable local minima, thereby delivering consistently superior solution quality.

### v2

@@ 7 Conclusion
We study a utility maximization problem under stochastic dominance constraints. Starting from an S-shaped utility function, we derive the explicit optimal solution without a liquidation boundary under first-order stochastic dominance (FSD) constraints. For the more challenging SSD problem, particularly under non-concavity, obtaining an explicit optimal solution is inherently difficult. Motivated by the structural properties of the analytical theorem in the concave case, we introduce a Poor-Performance-Region Algorithm (PPRA). This algorithm efficiently identifies the candidate structure that a potential optimal solution should satisfy and proves effective in the vast majority of cases. Extensive numerical experiments illustrate how the algorithm operates, confirm its broad applicability, and demonstrate its capability to handle the few exceptional cases. Building on the structural insights provided by the PPRA, we further recognize the potential of neural networks in tackling SSD problems under non-concavity. While neural networks offer strong approximation capabilities, their direct application often suffers from slow convergence and severe sensitivity to local optima induced by non-concavity. By leveraging the PPRA’s ability to capture the essential structure of the optimal solution, we develop an algorithm-guided piecewise-neural-network framework that integrates these structural cues into the learning process. Compared with a standard neural-network approach, this framework achieves significantly faster convergence and effectively avoids undesirable local minima, thereby delivering consistently superior solution quality.